OpenStudy (anonymous):

Medal and fan awarded :))) Given the following functions f(x) and g(x), solve (f + g)(3) and select the correct answer below: f(x) = 2x + 21 g(x) = x − 24 −27 −21 6 48

OpenStudy (anonymous):

@Mehek14

OpenStudy (welshfella):

first add f(x) and g(x)

OpenStudy (anonymous):

How do I add f(x) and g(x)? This is a brand new topic to me, lol

OpenStudy (welshfella):

add like terms 2x + 21 + x - 24

OpenStudy (welshfella):

- simple algebra

OpenStudy (anonymous):

3 x - 3

OpenStudy (welshfella):

right now all you need do is plug in x = 3 and work it out

OpenStudy (anonymous):

f(3) g(3)?

OpenStudy (welshfella):

3(3) - 3 = ?

OpenStudy (anonymous):

6

OpenStudy (welshfella):

right! thats it

OpenStudy (welshfella):

(f + g)(3) = 6

OpenStudy (anonymous):

So how do I work that out

OpenStudy (welshfella):

- you've worked it out 3(3) - 3 = 6

OpenStudy (welshfella):

(f + g)(3) means the sum of f(x) and g(x) when x = 3

OpenStudy (anonymous):

So the answer's 6?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

Great :D

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

Could you help me with a few more?

OpenStudy (welshfella):

well i've only got 5 minutes...

OpenStudy (welshfella):

ok i could help with 1 more

OpenStudy (anonymous):

Great :)

OpenStudy (anonymous):

Dishwashers are on sale for 25% off the original price (d), which can be expressed with the function p(d) = 0.75d. Local taxes are an additional 14% of the discounted price, which can be expressed with the function c(p) = 1.14p. Using this information, which of the following represents the final price of a dishwasher with the discount and taxes applied? c(p) ⋅ p(d) = 0.855pd c(p) + p(d) = 1.89d c[p(d)] = 0.855d d[c(p)] = 1.89p

OpenStudy (welshfella):

0.75d - 0.14*0.75d

OpenStudy (anonymous):

0.645?

OpenStudy (anonymous):

or is it .14

OpenStudy (welshfella):

- sorry gotta go right now

OpenStudy (anonymous):

Darn :(

OpenStudy (welshfella):

I havent got this clear in my mind yet

OpenStudy (anonymous):

Thanks for the help :/